### v1

# Hard rigid rods on Husimi trees

Abstract
We study the thermodynamic behavior of hard rigid rods of size $k$ (i.e., $k$-mers) on four- and six-coordinated Husimi lattices (HLs), respectively built with squares (square HL) and triangles (triangular HL). In both lattices, dimers ($k=2$) and trimers ($k=3$) only present a isotropic phase, whereas a isotropic-nematic transition is observed for $k\geq 4$. In the square HL, this transition is continuous and occurs at a critical monomer activity which displays a nonmonotonic variation with $k$, while the critical rod activity and density are always decreasing functions of $k$. The isotropic-nematic transition is discontinuous in the triangular HL, but the $k$-dependence of the coexistence activities and density is analogous to that found for the square case. No transition from the nematic to a high-density disordered phase is found in these HLs. In general, this scenario is very similar to that already observed for rods on the Bethe lattice, though the critical parameters obtained here are in most cases closer to those reported in the literature for the square and triangular lattices. The entropy per site of fully-packed rods is also investigated in detail in the triangular HL, where its value for dimers differs by only 0.7% from the exact result for the triangular lattice.
I Introduction
Particle systems with excluded volume interactions only have been studied for much time. Besides a disordered phase, they may also exhibit ordered phases and thus undergo entropy-driven transitions between them. Many different shapes have been considered for the particles, such as hard spheres, squares, triangles, tetraminoes, and so on cmc05 ; ps88 ; fal07 ; b80 ; vn99 ; bsg09 ; m05 . Some systems composed by binary binary and ternary mixtures ternary of hard particles have also been considered in the literature. In this paper, we investigate hard rigid rods, which are frequently associated to liquid crystals vl92 . The study of this particular system started with a seminal paper by Onsager o49 , where he concluded that, at sufficiently high densities and large aspect ratios, the rods would undergo a discontinuous phase transition between an isotropic and a nematic phase. While Onsager considered the rods in continuous space, Flory f56 placed them on a lattice, but still allowing their orientations to be continuous. He solved the problem in a mean-field approximation, finding also a transition between isotropic and nematic phases. Zwanzig z63 studied a system of hard rods in continuum space, but with discrete orientations, finding, as Onsager, a discontinuous isotropic-nematic phase transition.
Here, we are interested in rigid rods placed on lattices, formed by $k$ consecutive monomers aligned in one lattice direction, so that between consecutive monomers of a rod there is a lattice edge. We will call these rods $k$-mers. For the particular case of $k=2$ (dimers) on the square lattice, the entropy of the system in the limit of full coverage has been exactly determined, using both pfaffians k61 ; tf61 and transfer matrices l67 ; wp21 , being $s_{2}^{(sq)}=G/\pi=0.29156\ldots$ the entropy per site, where $G$ is Catalan’s constant. Fully-packed dimers on the triangular lattice have also a long history (see, e.g., Refs. Nagle66 ; Phares85 ; Kenyon00 ; Fendley02 ; Wu06 ), where $s_{2}^{(tri)}=0.42859\ldots$ Fendley02 ; Wu06 . The exact determination of the entropy for dimers in this limit is also possible for other two-dimensional lattices Wu06 , as well as in higher dimensions dc08 . For larger values of $k$, as far as we known, there are no exact results available for the full lattice entropy of $k$-mers placed on regular lattices. However, rather precise estimates have been obtained using transfer matrix methods for trimers gdj07 and other $k$-mers with $2\leq k\leq 10$ rsd23 on the square lattice, which agree with the outcomes from numerical simulations p21 . Semi-analytical solutions of fully-packed rods on generalized Husimi lattices — built with square lattice clusters of effective lateral size $L$ — also provide results in fair agreement with those for the square lattice, when extrapolated to the $L\rightarrow\infty$ limit rso22 . To the best of our knowledgement, there are no analogous results for the triangular lattice reported in the literature.
Turning our attention to the more general case where vacancies may exist between rods placed on two-dimensional lattices, we start mentioning that Monte Carlo simulations demonstrate that no nematic order exists for $k<k_{min}=7$, whereas a continuous isotropic-nematic transition is found for $k\geq 7$ gd07 . Its universality class depends on the number of possible orientations of the rods; and may be Ising mf08 ; mf08p ; fv09 or, more generally, Potts mf08 ; mf08p ; mf08pp . The existence of the nematic phase at intermediate densities was also rigorously established for large $k$’s ds13 . Interestingly, a second transition — from the nematic to a high-density disordered phase — is found at high rods’ densities gd07 , but its nature is not so well known. Although usual Monte Carlo algorithms using local evaporation and deposition moves are very inefficient in the high density region, a new procedure introduced some time ago (simultaneously updating all sites of a strip using transfer matrix procedures) reduced this difficulty k12 ; k13 , indicating a non-Ising transition for $k=7$. Nevertheless, recent results suggest this transition to be discontinuous sdr22 .
The thermodynamic behavior of $k$-mers was also studied on the Bethe lattice (BL) and Bethe-like lattices drs11 . There, a continuous isotropic-nematic transition was found for the four-coordinated lattice ($q=4$) already for $k\geq k_{min}^{(BL)}=4$; in contrast with $k_{min}=7$ obtained for the square lattice gd07 . For BLs with coordination $q\geq 6$, the isotropic-nematic transition is discontinuous and appears, again, for $k\geq k_{min}^{(BL)}=4$; which differs from the continuous transition found (for $k\geq 7$) in the triangular lattice mf08p . Furthermore, the transition from the nematic to the high density isotropic phase is missing in the BL for all rod size $k$ and coordination $q$. We recall that the exact (or numerically exact) solution of a given model on the BL corresponds to the result of the so-called mean-field Bethe approximation on a regular lattice with the same coordination number b82 . So, the differences above lead us to inquire whether the rods’ behavior can be better captured by more elaborated mean-field approaches. In order to address this, in this paper, we study the thermodynamic properties of $k$-mers, for general densities, on the four-coordinated Husimi lattice built with squares and on the six-coordinated Husimi lattice built with triangles. Since loops are present in these Husimi cacti, they are expected to provide better mean-field approximations for the models on the square and triangular lattices than the simpler BL solution h50 ; t76 . This was indeed the case in the full lattice limit analyzed in Ref. rso22 , as already mentioned above. As demonstrated in what follows, the scenario found here is not so different to that just described for the BL; even though, in most cases, we obtain critical parameters closer to those estimated in the literature for the regular lattices, provided by techniques such as Monte Carlo simulations or transfer matrix calculations.
The rest of the paper is organized as follows. In Sec. II the Husimi lattices are defined and the model is solved on the tree built with squares; while the solution on the triangular Husimi cactus is presented in the Appendix. The thermodynamic behavior of rods placed on the square and triangular Husimi lattices are shown and discussed in sections III and IV, respectively. Our final discussions may be found in Sec. V.
II Solution of the model on the Husimi lattice built with squares
The Husimi tree with coordination number $q=4$, as consider here, is just a treelike structure where successive generations of squares are added to a central square, as is shown in Fig 1(a). A similar tree, with coordination $q=6$, can be built up by joining generations of triangles to a central triangle, as done in Fig. 1(b). The solution of a given model in the core of these Husimi trees, in the thermodynamic limit, is called a Husimi lattice (HL) solution of this model in the literature t76 . The square HL (triangular HL) provides a mean-field approximation for the model’s behavior on the square (triangular) lattice.
Figure 1: Husimi trees built with (a) squares and (b) triangles, by starting with a central polygon and adding two generations of them at the surface sites of the previous generations. Following this procedure, such trees always have an elementary polygon [square in (a) and triangle in (b)] at their centers. The four-coordinated Cayley tree obtained by connecting the squares’ centers in (a) is represented by dashed lines.
The solution of the rigid rods model on the square HL will be presented here, while the one for the triangular case will be discussed in appendix A. The model we are analyzing is a generalization of the full lattice limit, which was recently considered by some of us on the square HL rso22 , but now empty sites are allowed and activities $z_{x}=\exp[\mu_{x}/(k_{B}T)]$ and $z_{y}=\exp[\mu_{y}/(k_{B}T)]$ are associated to each site occupied by a monomer in a rod aligned in the $x$ and $y$ direction, respectively. The quantities $\mu_{x}$ and $\mu_{y}$ are the chemical potentials of a monomer, which is part of a $k$-mer, so that we are in the grand-canonical ensemble. Moreover, $k_{B}$ is Boltzmann’s constant and $T$ is the temperature. Although we will restrict ourselves to the thermodynamic properties of the system in the case $z_{x}=z_{y}=z$, for latter convenience, the solution will be presented considering different activities. As expected, in all expressions below the full lattice limit results are recovered in the $z\to\infty$ limit.
As usual in solving models on such lattices, we start obtaining recursion relations for the grand-canonical partial partition functions (ppf’s) of rooted sub-trees for fixed configurations of their root site. In other words, we consider the operation of connecting three sub-trees to three corners of the root square of the sub-tree with an additional generation of squares. We will label the root configuration of the sub-tree according to the number of monomers already attached to the rod reaching it from above. If there is no such rod, this label will be equal to zero. If there is a reaching rod, the label will be the number of monomers already included in the rod with a second component indicating the direction of the rod, $x$ or $y$. Therefore, there are $12k-$ 1 ppf’s, whose first labels are in the range $[0,...,k-1]$. In Fig. 2 the possible sub-tree root configurations are shown.
Figure 2: Possible root configurations of a sub-tree. From the left to the right: no rod reaching on the root site from above ($g_{0}$), a rod in the $x$-direction arriving at the root site ($g_{(i,x)},\,i=1,2,...,k-1$), and a rod in the $y$-direction reaching on the root site ($g_{(i,y)},i=1,2,...,k-1$).
We now proceed obtaining the recursion relations for the ppf’s, by considering the operation of attaching three sub-trees to a new root square. The possible configurations of the new root square for the case of no rod arriving at the new root site are illustrated in Fig. 3.
Figure 3: Contributions to the recursion relation for $g^{\prime}_{0}$. The thick red lines indicate edges of a $k$-mer placed on the new root square.
Taking into account that the activity of a monomer on the root site of the ppf will be considered in the next iteration, the recursion relation for the partial partition function with no rod reaching on the root site coming from above is:
$g_{0}^{\prime}$ $=$ $[g_{0}+z_{x}g_{(k-1,x)}+z_{y}g_{(k-1,y)}]^{3}+[g_{0}+z_{x}g_{(k-1,x)}+z_{y}g_{(k-1,y)}]z_{x}^{2}\sum_{n=0}^{k-2}g_{(n,x)}g_{(k-2-n,x)}+$ (1)
$+[g_{0}+z_{x}g_{(k-1,x)}+z_{y}g_{(k-1,y)}]z_{y}^{2}\sum_{n=0}^{k-2}g_{(n,y)}g_{(k-2-n,y)},$
and the other recursion relations, for ppf’s of sub-trees with rods arriving at the root site, may be obtained using similar combinatorial arguments. The results are:
$g^{\prime}_{(i,x)}=z_{x}g_{(i-1,x)}\left\{[g_{0}+z_{x}g_{(k-1,x)}+z_{y}g_{(k-1,y)}]^{2}+z_{x}^{2}\sum_{n=0}^{k-2}g_{(n,x)}g_{(k-2-n,x)}\right\},$ (2)
and
$g^{\prime}_{(i,y)}=z_{y}g_{(i-1,y)}\left\{[g_{0}+z_{x}g_{(k-1,x)}+z_{y}g_{(k-1,y)}]^{2}+z_{y}^{2}\sum_{n=0}^{k-2}g_{(n,y)}g_{(k-2-n,y)}\right\},$ (3)
where $i=1,2,...,,k-1$. The thermodynamic limit of these ppf’s is attained when the number of iterations $N$ diverges, and, as expected, they also diverge in this limit, so it is convenient to define their ratios:
$R_{(i,x)}=\frac{g_{(i,x)}}{g_{0}},\quad\text{ and }\quad R_{(i,y)}=\frac{g_{(i,y)}}{g_{0}}$ (4)
for $i=1,2,...,k-1$. We notice that, if we write the recursion relations for the ppf’s as $g^{\prime}_{(i,x)}=g_{0}^{3}f_{(i,x)}$ and $g^{\prime}_{(i,y)}=g_{0}^{3}f_{(i,y)}$, the functions $f_{i}$ depend only of the activities and the ratios defined above. It is easy to obtain the $2(k-1)$ recursion relations for the ratios. They are:
$R^{\prime}_{(i,x)}=z_{x}R_{(i-1,x)}\frac{A^{2}+S_{x}}{A^{3}+A(S_{x}+S_{y})},$ (5)
and
$R^{\prime}_{(i,y)}=z_{y}R_{(i-1,y)}\frac{A^{2}+S_{y}}{A^{3}+A(S_{x}+S_{y})},$ (6)
where $A=1+z_{x}R_{(k-1,x)}+z_{y}R_{(k-1,y)}$, with $i=1,2,...,k-1$ and
$S_{x}$ $=$ $z_{x}^{2}\sum_{n=0}^{k-2}R_{(n,x)}R_{(k-2-n,x)},$
$S_{y}$ $=$ $z_{y}^{2}\sum_{n=0}^{k-2}R_{(n,y)}R_{(k-2-n,y)},$ (7)
noticing that, from the definition, $R_{(0,x)}=R_{(0,y)}=1$. The recursion relations suggest the following ansatz for the fixed point values of the ratios, attained by them in the thermodynamic limit:
$R_{(i,x)}^{*}=x_{1}^{i}\quad\text{ and }\quad R_{(i,y)}^{*}=x_{2}^{i}.$ (8)
Substitution of this ansatz in the fixed point equation $R_{(i,x)}^{\prime}=R_{(i,x)}$ and $R_{(i,y)}^{\prime}=R_{(i,y)}$ leads to the following equations:
$x_{1}[(1+H_{k-1})^{3}+(k-1)(1+H_{k-1})H_{k-2}]$ $=$ $z_{x}[(1+H_{k-1})^{2}+(k-1)z_{x}^{2}x_{1}^{k-2}]$ (9a)
$x_{2}[(1+H_{k-1})^{3}+(k-1)(1+H_{k-1})H_{k-2}]$ $=$ $z_{y}[(1+H_{k-1})^{2}+(k-1)z_{y}^{2}x_{2}^{k-2}],$ (9b)
where $H_{k-i}=z_{x}^{i}x_{1}^{k-i}+z_{y}^{i}x_{2}^{k-i}$. By either (numerically) solving these equations or iterating the recursion relations (Eqs. 5 and 6), for given $z_{x}=z_{y}=z$ and $k$, we obtain the fixed points which define the thermodynamic properties of the phases of the model on the square HL.
The grand-canonical partition function of the system may be obtained considering the operation of connecting four sub-trees with $N$ generations to the central square of the Husimi tree. A convenient alternative is to connect a sub-tree with $N+1$ generations to another one with $N$ generations, so that they share their root sites. This common root site may be empty or occupied by a monomer of a horizontal or vertical rod, thus:
$Y_{N}=g_{0}^{\prime}g_{0}+\sum_{n=0}^{k-1}(z_{x}g^{\prime}_{(n,x)}g_{(k-n-1,x)}+z_{y}g^{\prime}_{(n,y)}g_{(k-n-1,y)})=g_{0}^{4}y,$ (10)
with
$y$ $=$ $f_{0}+\sum_{n=0}^{k-1}(z_{x}f_{(n,x)}R_{(k-n-1,x)}+z_{y}f_{(n,y)}R_{(k-n-1,y)})=$ (11)
$f_{0}\left[1+\sum_{n=0}^{k-1}(z_{x}R^{\prime}_{(n,x)}R_{(k-n-1,x)}+z_{y}R^{\prime}_{(n,y)}R_{(k-n-1,y)})\right],$
where
$f_{0}=A^{3}+A(S_{x}+S_{y}),$ (12)
and it is implicit that $g_{0}$ corresponds to the ppf of a sub-tree with an empty root site and $N$ generations. The average number of $k$-mers in the horizontal and vertical directions reaching the site of the central square of the tree where the two sub-trees meet are, respectively:
$n_{x}$ $=$ $\frac{z_{x}\sum_{n=0}^{k-1}R^{\prime}_{(n,x)}R_{(k-n-1,x)}}{1+\sum_{n=0}^{k-1}(z_{x}R^{\prime}_{(n,x)}R_{(k-n-1,x)}+z_{y}R^{\prime}_{(n,y)}R_{(k-n-1,y)})},$ (13a)
$n_{y}$ $=$ $\frac{z_{y}\sum_{n=0}^{k-1}R^{\prime}_{(n,y)}R_{(k-n-1,y)}}{1+\sum_{n=0}^{k-1}(z_{x}R^{\prime}_{(n,x)}R_{(k-n-1,x)}+z_{y}R^{\prime}_{(n,y)}R_{(k-n-1,y)})},$ (13b)
so that the mean density of sites occupied by monomers is
$\rho=n_{x}+n_{y},$ (14)
and a nematic order parameter may be defined as
$\Psi=\frac{|n_{x}-n_{y}|}{n_{x}+n_{y}}.$ (15)
The free energy of the system on the whole Husimi tree, in the thermodynamic limit, is dominated by the surface. The bulk grand-canonical free energy per site, limited to the central region, discarding the contribution of the surface, is then considered Gujrati ; MinosJurgen ; tiagoPol . Basically, one assumes that the free energy per site of sites located on the surface of the tree is $\phi_{s}$ and for sites in the bulk it is equal to $\phi_{b}$. The result for the bulk free energy per site is:
$\phi_{b}=-\frac{k_{B}T}{4}\ln\left[\frac{Y^{\prime}}{Y^{3}}\right],$ (16)
where $Y$ and $Y^{\prime}$ are the partition functions of Husimi trees with $N$ and $N+1$ generations, respectively. It may then be shown that
$\varphi_{b}\equiv\frac{\phi_{b}}{k_{B}T}=-\ln\left[\frac{f_{0}}{y^{1/2}}\right].$ (17)
In the thermodynamic limit ($N\to\infty$) the ratios should be replaced by their fixed point values $R_{(i,x)}^{*}=x_{1}^{i}$ and $R_{(i,y)}^{*}=x_{2}^{i}$, where $x_{1}$ and $x_{2}$ are the solutions of the fixed point equations (9).
In the calculations above we defined the Husimi tree so that the center of an elementary square is located at at its center. We could, alternatively, define it in such a way that a site of the tree is located at the center. In this case, although we would expect different results for a finite tree, the average values at the central region of both trees will be equal in the thermodynamic limit. The partition function in this case is obtained connecting two sub-trees with the same number of generations to the central site, giving
$Y_{a,N}=g_{0}^{2}+\sum_{n=0}^{k-1}[z_{x}g_{(n,x)}g_{(k-n-1,x)}+z_{y}g_{(n,y)}g_{(k-n-1,y)}]=g_{o}^{2}y_{a},$ (18)
where now
$y_{a}=1+\sum_{n=0}^{k-1}[z_{x}R_{(n,x)}R_{(k-n-1,x)}+z_{y}R_{(n,y)}R_{(k-n-1,y)}].$ (19)
The average number of $k$-mers in the horizontal and vertical direction reaching the central site in the thermodynamic limit may be expressed in terms of the parameters $x_{1}$ and $x_{2}$ (Eq. 8) as
$n_{x}=$ $=$ $\frac{z_{x}kx_{1}^{k-1}}{1+k(z_{x}x_{1}^{k-1}+z_{y}x_{2}^{k-1})},$ (20a)
$n_{y}=$ $=$ $\frac{z_{y}kx_{2}^{k-1}}{1+k(z_{x}x_{1}^{k-1}+z_{y}x_{2}^{k-1})}.$ (20b)
and a procedure similar to the one used above for the square-centered tree leads to the bulk free energy per site:
$\varphi_{b}\equiv\frac{\phi_{b}}{k_{B}T}=-\frac{1}{2}\ln\left[\frac{f_{0}}{y_{a}}\right],$ (21)
where $f_{0}$ is given in Eq. 12
The entropy per site may be obtained considering that $\phi_{b}=u_{b}-Ts-\rho\mu$, so that
$\frac{s}{k_{B}}=-\varphi_{b}-\rho\ln z,$ (22)
for $u_{b}=0$, since the system is athermal.
Finally, to study the stability of the fixed points, we obtained the $(2k-2)\times(2k-2)$ Jacobian matrix of the recursion relations of the ratios of ppf’s, whose elements are:
$J_{i,j}=\left(\frac{\partial R^{\prime}_{i}}{R_{j}}\right).$ (23)
A particular fixed point will be stable if the modulus of the leading eigenvalue of the Jacobian is smaller than one.
III Results for rods on the square HL
To obtain the thermodynamic properties of the model on the Husimi lattice, we start, for fixed values of the monomer activity $z$ and number of monomers in the rod $k$, calculating the fixed points of the recursion relations in Eqs. 5 and 6. This may be done by iterating such recursion relations, but it is simpler to solve the fixed point equations 9 and the ratios may then be obtained through expressions 8. For $k=2,3$, we find that only the isotropic fixed point exists ($x_{1}=x_{2}=x)$, for any value of the activity $z$, so that the nematic order parameter vanishes identically. For $k\geq 4$ an isotropic fixed point is stable for small activities, while for higher activities a nematic fixed point is stable. The stability limits of both fixed points are coincident [at a critical activity $z_{c}(k)$], indicating that the transition between the isotropic and the nematic phase is continuous.
This is indeed confirmed in Fig. 4, where the order parameter is shown as a function of the density of sites occupied by the rods on the lattice, for different values of $k$. As $k$ grows, $\Psi$ approaches unity in the limit of full occupancy (corresponding to $z\rightarrow\infty$ and, thus, $\rho=1$), but it does not reach it, as already demonstrated in Ref. rso22 . Therefore, in such a limit the stable phase is nematic for all $k\geq 4$ and no nematic-to-isotropic transition exists at high densities, similarly to the BL solution drs11 .
Figure 4: Nematic order parameter as a function of the density of sites occupied by monomers, for $k$-mers placed on the square HL. From right to left, results for $k=4,5,6,$ and $7$.
The values of the critical activities, densities, and entropies for $4\leq k\leq 15$, are summarized in Table 1. We notice that the critical densities and entropies are monotonically decreasing functions of $k$, but the critical activities show a minimum at $k=10$, followed by a slow increase. We may compare the critical parameters with the ones found for the model on the four coordinated Bethe lattice in drs11 , which are:
$z_{c}^{(BL)}=\frac{(k-1)^{\frac{2k-2}{k}}}{k(k-3)},$ (24)
$\rho_{c}^{(BL)}=\frac{2}{k-1},$ (25)
and
$s_{c}^{(BL)}$ $=$ $2\left(1-\frac{1}{k}\right)\ln\left(1-\frac{1}{k}\right)-\left(1-\frac{2}{k-1}\right)\ln\left(1-\frac{2}{k-1}\right)-$ (26)
$2\left(\frac{1}{k(k-1)}\right)\ln\left(\frac{1}{k(k-1)}\right).$
We found that, as expected, the absolute values of the relative differences between the values of these parameters on the two lattices are decreasing functions of $k$, starting with values of the order of 20% for $k=4$ and ending with values around 0.5% or lower.
Table 1: Critical activity ($z_{c}$), density ($\rho_{c}$) and entropy ($s_{c}$) for rods with different number of monomers $k$ placed on the square HL.
$k$ $z_{c}$ $\rho_{c}$ $s_{c}$
4 1.675089 0.761905 0.320403
5 0.964988 0.535714 0.293498
6 0.827138 0.417391 0.233700
7 0.777397 0.343137 0.186956
8 0.756853 0.291793 0.152346
9 0.748910 0.254032 0.126464
10 0.747170 0.225035 0.106714
11 0.748709 0.202040 0.091330
12 0.752040 0.183346 0.079117
13 0.756352 0.167840 0.069258
14 0.761179 0.154767 0.061181
15 0.766244 0.143594 0.054477
The entropy per lattice site, also as a function of the density of sites occupied by monomers, is depicted in Fig. 5. It obviously starts at zero for $\rho=0$, reaches a maximum at an intermediate density, which grows monotonically with $k$, and ends at the values already obtained in Ref. rso22 in the full lattice limit.
Figure 5: Entropy per lattice site (in units of $k_{B}$) as a function of the density of sites occupied by monomers, for $k$-mers placed on the square HL. Downwards: results for $k=2,3,4,$ and $7$.
As expected, when we compare the results found on the Husimi lattice with the behavior of the model on the Bethe lattice drs11 , in general, the results on the Husimi lattice are somewhat closer to the estimates provided by numerical simulations of the system on the square lattice. For instance, as discussed in the Introduction, on the square lattice the nematic order parameter vanishes in the full lattice limit, while both in the Bethe and Husimi lattices it does not vanish. However, as demonstrated in rso22 , $\Psi$ is always smaller on the Husimi lattice than on the Bethe lattice; see table III of Ref. rso22 . It is also noteworthy in such a table, as it was seen in the comparisons between critical parameters on the Bethe and Husimi lattices above, that the difference between the order parameters on the two lattices becomes smaller as $k$ increases, since the ratio between the size of the rods and the size of the elementary squares of the lattice grows. This effect is observed in all comparisons between both lattices, being clearly visible, e.g., in the full lattice entropies presented in table III of Ref. rso22 .
The values of the critical density on the HL as a function of the size of the rod $k$ are compared with those for the BL and for the square lattice in Fig. 6. The latter ones were obtained by Matoz-Fernandez et al. mf08 ; mf08p , according to which $\rho_{c}(k)\sim k^{-1}$ and, moreover, $\rho_{c}(10)\approx 0.502$ for rigid rods on the square lattice, such that $\rho_{c}(k)\simeq 5.02/k$. We notice that, for $k=7$ this expression gives $\rho_{c}(7)\approx 0.717$, which is not so different from the estimate reported by Kundu et al. k13 : $\rho_{c}(7)\approx 0.745$. Recalling that results on the treelike lattices are mean-field approximations of the model on the square lattice, we expect them to underestimate the critical density. This is indeed confirmed in Fig. 6, where we may see that the critical densities on the square HL are larger than those found on the Bethe lattice. Of course, the mean-field results are still much lower than the available estimates for the square lattice ones.
Figure 6: Critical densities on Bethe and square HL, as functions of $k$. Circles correspond to the four coordinated Bethe lattice results drs11 , squares to Husimi lattice results and the blue line are simulational estimates for the model on the square lattice mf08 ; mf08p
IV Results for rods on the triangular HL
IV.1 Isotropic-nematic transition
Similarly to the results above for the square HL (and also to those for the BL drs11 ), the $k$-mers never present nematic order in the triangular HL for $k=2$ and $3$, whereas an isotropic-nematic transition is found for $k\geq 4$. Such a transition is always discontinuous in the triangular HL case, in consonance with the BL approach for coordination $q=6$ drs11 . This is shown in Fig. 7, where one sees that both the monomer density $\rho$ and the order parameter $\Psi$ present a discontinuous behavior as $z$ increases.
Figure 7: Results for rigid rods placed on the triangular HL. (a) Monomer density $\rho$ versus activity $z$, for rods with $k=5$, in the isotropic ($I$) and nematic ($N$) phases. (b) Nematic order parameter $\Psi$ as a function of $z$, for the indicated $k$’s. Similar results are found for all $k\geq 4$.
The coexistence points $z^{*}$ — estimated from the condition $\phi_{b}^{(I)}=\phi_{b}^{(N)}$, for a given $k$ — are summarized in Table 2 and displayed also in Fig. 8(a). As was found for the critical activities on the square HL, $z^{*}$ displays a non-monotonic variation with $k$: it decreases fast for small rods; then, presents a minimum at $k=12$; after which it passes to grow very slowly. On the other hand, the monomer densities at the coexistence, $\rho^{*(I)}$ and $\rho^{*(N)}$, are always decreasing functions of $k$ (at least for $k\leq 20$, which was the largest case analyzed here), as seen in Tab. 2 and Figs. 8(c) and 8(d). The gap between these densities $\Delta\rho^{*}=\rho^{*(N)}-\rho^{*(I)}$ is small, being $\Delta\rho^{*}\lesssim 0.06$, though the relative difference $\Delta\rho^{*}/\rho^{*(N)}$ increases monotonously with $k$, since $\rho^{*(N)}$ decreases faster than $\Delta\rho^{*}$.
As may be observed in Fig. 7(b), in Table 2 and in Fig. 8(b), the jump in the order parameter at coexistence, $\Psi^{*}$, increases with $k$, suggesting that the discontinuous character of the transition gets stronger for larger $k$’s. In contrast, however, the coexistence region decreases with $k$. Indeed, the activities, $z_{s}$, at the spinodals (estimated from the largest eigenvalue of Jacobian $|\Lambda|=1$) of the isotropic and nematic phases are also displayed in Tab. 2, where one observes that $\Delta z_{s}=z_{s}^{(I)}-z_{s}^{(N)}$ decreases (very slowly asymptotically) with the rods’ size.
Figure 8: Comparison between results for $k$-mers placed on the triangular HL and on the BL, at the coexistence points: (a) activity $z^{*}$, (b) monomer density of the nematic phase $\rho^{*(N)}$ and (c) nematic order parameter $\Psi^{*}$ versus $k$. The insertion in panel (a) highlights the behavior of $z^{*}$ for large $k$.
Noteworthy, the scenario above is very similar to that found for rods on the BL with $q=6$ drs11 . It is important to remark, however, that the coexistence points were estimated in Ref. drs11 by writing down the entropy [Eq. (32) there] as a function of $\rho$ and $\Psi$, and then defining the “critical density” $\rho_{c}$ as the value for which $s(\Psi,\rho=\rho_{c})$ has two peaks of equal height  [^1] When comparing the results here with those in Ref. drs11 , it is important to bear in mind that the activity there (let us call it $\tilde{z}$) is for rod, while here $z$ is for a single monomer, such that $z=\tilde{z}^{1/k}$.. We have revisited such calculations for the six-coordinate BL here, but determining the coexistence points through the free energy — i.e., by requiring that $\phi_{b}^{(I)}=\phi_{b}^{(N)}$ —, which gives results slightly different from those reported in drs11 .
Table 2: Coexistence activity ($z^{*}$), densities ($\rho^{*(I)}$ and $\rho^{*(N)}$) and order parameter ($\Psi^{*}$), and stability limits ($z_{s}^{(I)}$ and $z_{s}^{(N)}$) for rods of different sizes $k$ placed on the triangular HL.
$k$ $z^{*}$ $\rho^{*(I)}$ $\rho^{*(N)}$ $\Psi^{*}$ $z_{s}^{(I)}$ $z_{s}^{(N)}$
4 2.049172 0.773600 0.793726 0.517008 2.347627 2.015605
5 1.170803 0.592492 0.634234 0.640842 1.285300 1.157497
6 0.959370 0.479949 0.534063 0.709700 1.035761 0.950479
7 0.874695 0.403266 0.463657 0.752365 0.934285 0.867787
8 0.834113 0.347676 0.410778 0.780991 0.883966 0.828363
9 0.813224 0.305538 0.369300 0.801365 0.856576 0.808250
10 0.802429 0.272499 0.335746 0.816533 0.841056 0.798020
11 0.797270 0.245902 0.307967 0.828226 0.832262 0.793294
12 0.795437 0.224031 0.284549 0.837496 0.827522 0.791806
13 0.795638 0.205731 0.264514 0.845015 0.825328 0.792290
14 0.797109 0.190193 0.247165 0.851228 0.824783 0.793999
15 0.799378 0.176836 0.231985 0.856445 0.825325 0.796471
Figures 8(a)-(d) present the outcomes from such BL solution, comparing them with those for the triangular HL. In general, all quantities have larger values on the BL, but the same qualitative behavior. For instance, $z^{*}$ has also a non-monotonic variation with $k$ (with a minimum at $k=13$) in the BL case. The same non-monotonic behavior is observed in the spinodals of both the triangular HL (see Tab. 2) and the BL. In fact, for the BL the spinodal of the isotropic phase can be exactly determined, being drs11
$z_{s}^{(I,BL)}=\frac{\left(1-\frac{1}{k}\right)^{\frac{k-1}{k}}}{\left(1-\frac{3}{k-1}\right)[k(k-1)]^{\frac{1}{z}}},$ (27)
which displays a minimum at $k=16$, considering that $k\in\mathbb{Z}$. It is interesting that $z_{s}^{(I,BL)}\rightarrow\infty$ for $k=4$, meaning that the isotropic phase is always stable (or metastable) in the BL. In the triangular HL, on the other hand, $z_{s}^{(I)}$ is finite (and small) for all $k\geq 4$ (see Tab. 2). Actually, for all $k\geq 4$, we find that the size of the coexistence region $\Delta z_{s}=z_{s}^{(I)}-z_{s}^{(N)}$ is larger in the BL case.
Moreover, although the transitions are discontinuous in both the BL and triangular HL solutions, the nematic order parameters at coexistence are much smaller in the HL case, especially for the smaller $k$’s. These results demonstrate that, by improving the approximation, one obtains a behavior closer to a continuous transition, as expected for the triangular lattice mf08 ; mf08p . Curiously, however, the coexistence densities do not follow this trend. In fact, from Monte Carlo simulations of rigid rods on the triangular lattice and some theoretical approaches, Matoz-Fernandez et al. mf08 ; mf08p found that the continuous isotropic-nematic transitions occur at critical densities given by $\rho_{c}(k)\simeq 5.31/k$, for $k\geq 7$. As it may be observed in Figs. 8(c) and 8(d), these critical densities are much larger than the coexistence ones found here. Moreover, both $\rho^{*(I)}(k)$ and $\rho^{*(N)}(k)$, are more distant from $\rho_{c}(k)$ in the HL case.
IV.2 The full-packing limit
Now, we will discuss the case where all sites of the triangular HL are occupied by rods, corresponding to the $z\rightarrow\infty$ limit (and $\rho=1$) in the solution presented in the appendix.
The entropies of the isotropic and nematic phases are presented in Table 3, for $2\leq k\leq 7$. In agreement with the results from the previous subsection, only the isotropic phase is stable in the full lattice case for dimers and trimmers. For larger $k$’s, however, the entropy of the isotropic phase is negative, while that for the nematic phase is positive. Therefore, for $k\geq 4$, the nematic phase is the single one stable for large $z$ and this persists until the fully occupancy limit. Hence, no nematic-isotropic transition close to full-packing — as observed for rigid rods on the square lattice k12 — occurs in the triangular HL. The nematic order parameters are also shown in Tab. 3, which are close, but smaller than 1.
For comparison, Table 3 presents also the entropies and order parameters for rigid rods placed on the Bethe lattice (BL) with coordination $q=6$, in the full lattice limit, as obtained in Ref. drs11 . As in the HL case, the stable phase is isotropic for $k\leq 3$ and nematic for $k\geq 4$. Interestingly, the entropies of these phases are very similar in the BL and HL approximations. For instance, they differ by less than 2% in the case of dimers, for which they are very close also to the exact value on the triangular lattice: $s(3)=0.428594537\ldots$ (see, e.g., Ref. Wu06 and references therein). Remarkably, the difference between the HL and the exact result for dimers is only 0.7%. This is certainly a consequence of the fact that the dimer-dimer correlations on the triangular lattice are short ranged Fendley02 , so that the weakening of correlations caused by the treelike structure of the HL has a small effect on the properties of this model. Also, this suggests that generalized triangular HLs — built, e.g., with hexagonal triangular lattice cells, in the same spirit of the generalized square HLs considered in Refs. ro21 ; rso22 — may give the exact solution of the dimer problem for finite (and perhaps small) cells.
Table 3: Entropy $s$ of fully-packed rods placed on the Bethe lattice (BL) with coordination $q=6$ and on the triangular HL, for the isotropic ($I$) and nematic ($N$) phases, and several $k$’s. The values of the nematic order parameter $\Psi$ are also shown into parentheses.
$k$ 2 3 4 5 6 7
$s_{BL}^{(I)}$ 0.4400758 0.1460079 -2.605800E-02 -0.1407308 -0.2233532 -0.2860802
$s_{BL}^{(N)}$ 9.362827E-03 6.578885E-04 4.301164E-05 2.429306E-06
$(\Psi)$ (0.930306) (0.994925) (0.999612) (0.999974)
$s_{HL}^{(I)}$ 0.4315231 0.1189678 -7.192051E-02 -0.2034984 -0.3008850 -0.3764640
$s_{HL}^{(N)}$ 8.668699E-03 6.507839E-04 4.295364E-05 2.428988E-06
$(\Psi)$ (0.941627) (0.995035) (0.999612) (0.999974)
As demonstrated in Table III of Ref. rso22 , which shows the same results as presented in Table 3 here, but for the BL with $q=4$ and for the square HL — the rods’ entropies in these four-coordinated lattices are smaller than those in the six-coordinated cases analyzed here, as expected. Moreover, while the entropy for a given $k$ converges from below in $q=4$ case (see Tabs. III and V, and Fig. 2 in Ref. rso22 ), for the lattices considered here (with $q=6$) it converges from above to the asymptotic (triangular lattice) behavior. Namely, the number of possible $k$-mers’ configurations decreases as one goes from the BL ($q=6$) to the triangular HL and, then, to the triangular lattice; while the opposite behavior is observed for the lattices with $q=4$.
V Discussion
We have studied the model of $k$-mers, rigid rods composed by $k$ monomers placed on lattice sites with $k-1$ edges of the lattice between them, all aligned in the same direction, with excluded volume interactions only; that is, only configurations of the rods where each lattice site is occupied by at most one monomer are accepted. We obtained the solution of this model in the grand-canonical ensemble in the central region of two treelike lattices built with polygons (Husimi trees): the tree built with squares, so that two squares meet at each lattice site (ramification of squares equal to one and coordination number $q=4$); and the one built with triangles, with three of them meeting at each site (ramification of triangles equal to two and $q=6$). The exact solution in the core of these trees corresponds to a mean-field approximation of the model on the square and triangular lattice, respectively. Due to the existence of loops in the Husime trees, these calculations may be considered to be higher order mean-field approximations to the problem as compared to the Bethe lattice solution, which was obtained in Ref. drs11 .
On the Husimi lattice built with squares, the rods display a continuous isotropic-nematic transition, as it is also the case on the four-coordinated Bethe lattice and on numerical simulations on the square lattice. In contrast, a discontinuous transition was found for the model on the Husimi lattice built with triangles, as happens also on the six-coordinated Bethe lattice, while numerical simulations on the triangular lattice suggest a continuous transition in the class of the 3-state Potts model. Therefore, in this case, the Husimi lattice solution also does not lead to the order of the transition obtained using numerical simulations. It is worthy remarking, however, that mean-field approaches to the 3-state Potts model are known to display a discontinuous transition w82 , which suggests that the same thing can be happen in the rod model. Namely, the discontinuous transition found for $k$-mers in the mean-field solution on the triangular Husimi lattice may be viewed as an indication that this transition is indeed in the 3-state Potts class in the triangular lattice. We notice, for instance, that a similar scenario was recently found for an associating lattice gas model, which exhibits a continuous transition in the 3-state Potts class between a fluid a high-density liquid phase, as revealed by Monte Carlo simulations and transfer matrix calculations furlanALG ; ibagon21 , but it was found to be discontinuous in Husimi lattice solutions tiago10 ; furlanALG . Furthermore, we recall that numerical simulations of the rod model on the cubic lattice vdr17 suggested the isotropic-nematic transition to be continuous for $k\geq 7$, but the possibility of a weak discontinuous first-order transition, as is found for the 3-state Potts model on this lattice w82 , was not ruled out, although it is not clearly supported by the simulational results. Once again, the isotropic-nematic transitions found in the early studies via mean-field approximations of the three-dimensional models mentioned before o49 ; f56 ; z63 were also discontinuous.
If we consider the critical densities at which the transition happen in our calculations for the square case, they are always between the ones found on the Bethe lattice and the estimates coming from the simulations. The same is true for the entropies in the full lattice limit, so that for these parameters the Husimi lattice results are closer to the ones on the corresponding Bravais lattices than the ones coming from the solution on the Bethe lattice, as expected. On the other hand, the coexistence densities obtained in the triangular case display an opposite trend, suggesting that they converge non-monotonically to the triangular lattice results as the mean-field approximation is improved. Another point where there was no improvement in the Husimi lattice solutions when compared to the ones on the Bethe lattice was the minimum value of $k_{min}$ for which the phase transition was observed, which is equal to four in all cases, as opposed to $k_{min}=7$ for the square and triangular lattices. Calculations in the full lattice limit rso22 on treelike lattices built with larger square cells, such as the Kobayashi-Susuki lattice ks93 , indicate that larger values of $k_{min}$ are found when the size of the cells grows. Therefore, it is interesting to extend this study to include vacancies in the configurations of the model on such generalized Husimi lattices to find out the complete thermodynamic behavior of rigid rods on them. In a similar vein, investigating the $k$-mers on such lattices built with triangular lattice cells can be also very important to verify whether the isotropic-nematic transition turns out to be continuous at some point, as well as to estimate the entropies in the full-packing limit, since it seems that they have never been calculated in the literature for the triangular lattice, beyond the dimer case. We are currently working in these projects.
ACKNOWLEDGMENTS
This work used computational resources of the “Centro Nacional de Processamento de Alto Desempenho” in São Paulo (CENAPAD-SP, FAPESP). The authors acknowledge financial support by FAPERJ, FAPEMIG, CAPES and CNPq (Brazilian agencies).
Appendix A Solution of the model on the Husimi lattice built with triangles
We may obtain the solution of the rod problem on the triangular HL [as defined in Fig. 1(b)] following the same lines as in the square HL case discussed in Sec. II. The main differences are that now we have three lattice directions — which will be labeled as $A$, $B$ and $C$ here — and two sub-trees are attached to each of the two non-rooted sites of a rooted triangle, in order to obtain a lattice with coordination $q=6$; see Figs. 1(b) and 9. Since there are three lattice directions, but only two edges arriving at a root site, at first, one should distinguish among the situations where such edges are of type $AB$, $BC$ or $CA$. To keep the notation as simple as possible, we will name the ppf’s according to the label of the direction opposite (i.e., not connecting) to the root site. Thereby, for the configurations with no reaching rod (from above) in the root site, we will have the ppf’s $a_{0}$, $b_{0}$ and $c_{0}$. In the same way, we shall have the ppf’s $b_{(i,A)}$ and $c_{(i,A)}$; $a_{(i,B)}$ and $c_{(i,B)}$; and $a_{(i,C)}$ and $b_{(i,C)}$, when the reaching rod has $i$ monomers incorporated to it in direction $A$, $B$ or $C$. By defining the sums:
$s_{A}=b_{0}c_{0}+z_{A}\sum_{n=0}^{k-1}b_{(n,A)}c_{(k-1-n,A)}+z_{B}b_{0}c_{(k-1,B)}+z_{C}c_{0}b_{(k-1,C)},$ (28a)
$s_{B}=a_{0}c_{0}+z_{A}a_{0}c_{(k-1,A)}+z_{B}\sum_{n=0}^{k-1}a_{(n,B)}c_{(k-1-n,B)}+z_{C}c_{0}a_{(k-1,C)},$ (28b)
$s_{C}=a_{0}b_{0}+z_{A}a_{0}b_{(k-1,A)}+z_{B}b_{0}a_{(k-1,B)}+z_{C}\sum_{n=0}^{k-1}a_{(k-1-n,C)}b_{(n,C)},$ (28c)
the recursion relations (RRs) for the ppf’s related to the opposite direction $A$ are given by
$a^{\prime}_{0}$ $=$ $s_{B}s_{C}+z_{A}^{2}a_{0}^{2}\sum_{n=0}^{k-2}b_{(n,A)}c_{(k-2-n,A)},$ (29a)
$a^{\prime}_{(i,B)}$ $=$ $z_{B}b_{0}a_{(i-1,B)}s_{B},$ (29b)
$a^{\prime}_{(i,C)}$ $=$ $z_{C}c_{0}a_{(i-1,C)}s_{C}.$ (29c)
where $i=1,\ldots,k-1$, for $k\geqslant 2$. The analogous RRs for $b$ and $c$ can be obtained from these expressions through cyclic permutations of type $a\rightarrow b$, $b\rightarrow c$ and $c\rightarrow a$, along with $A\rightarrow B$, $B\rightarrow C$ and $C\rightarrow A$.
Figure 9: Possible states of the root sites for rods placed on the triangular HL. The thicker (red) lines represent a reaching rod with $i$ monomers incorporated to it. The lattice directions ($A$, $B$ and $C$) are also indicated.
Then, by defining the ratios $F_{(i,B)}=a_{(i,B)}/a_{0}$, $F_{(i,C)}=a_{(i,C)}/a_{0}$, $G_{(i,A)}=b_{(i,A)}/b_{0}$, $G_{(i,C)}=b_{(i,C)}/b_{0}$, $H_{(i,A)}=c_{(i,A)}/c_{0}$ and $H_{(i,B)}=c_{(i,B)}/c_{0}$, one readily obtains a set of $(6k-6)$ RRs for them, whose fixed points define the phases of the system. It turns out, however, that such RRs always converge to fixed points where $G_{(i,A)}=H_{(i,A)}$, $F_{(i,B)}=H_{(i,B)}$ and $F_{(i,C)}=G_{(i,C)}$, demonstrating that there is no need to distinguish between the different directions of the opposite edges. Therefore, we can simplify the solution, by considering a single ppf $g_{0}$ (instead of $a_{0}$, $b_{0}$ and $c_{0}$) when no rod arrives at the root site (from above), and ppf’s $g_{(i,j)}$ if a rod with $i=1,\ldots,k-1$ monomers incorporated reaches the root site in direction $j=A,B$ or $C$. See Fig. 9 for a illustration of these definitions. In this way, the number of ppf’s reduces to only $23k-$ 2, with the related RRs given by
$g^{\prime}_{0}$ $=$ $s_{A}s_{C}+z_{B}^{2}g_{0}^{2}\sum_{n=0}^{k-2}g_{(n,B)}g_{(k-n-2,B)},$ (30a)
$g^{\prime}_{(i,A)}$ $=$ $z_{A}g_{0}g_{(i-1,A)}s_{A},$ (30b)
$g^{\prime}_{(i,B)}$ $=$ $z_{B}g_{0}g_{(i-1,B)}s_{B},$ (30c)
$g^{\prime}_{(i,C)}$ $=$ $z_{C}g_{0}g_{(i-1,C)}s_{C},$ (30d)
where the sums defined in Eqs. 28 simplify to
$s_{A}=g_{0}^{2}+z_{A}\sum_{n=0}^{k-1}g_{(n,A)}g_{(k-n-1,A)}+z_{B}g_{0}g_{(k-1,B)}+z_{C}g_{0}g_{(k-1,C)},$ (31a)
$s_{B}=g_{0}^{2}+z_{A}g_{0}g_{(k-1,A)}+z_{B}\sum_{n=0}^{k-1}g_{(n,B)}g_{(k-n-1,B)}+z_{C}g_{0}g_{(k-1,C)},$ (31b)
$s_{C}=g_{0}^{2}+z_{A}g_{0}g_{(k-1,A)}+z_{B}g_{0}g_{(k-1,B)}+z_{C}\sum_{n=0}^{k-1}g_{(n,C)}g_{(k-n-1,C)}.$ (31c)
Note that $g_{(0,A)}=g_{(0,B)}=g_{(0,C)}=g_{0}$ in these equations. Then, we may defined the (new) ratios of ppf’s as:
$A_{i}=\frac{g_{(i,A)}}{g_{0}},\quad B_{i}=\frac{g_{(i,B)}}{g_{0}},\quad\text{and}\quad C_{i}=\frac{g_{(i,C)}}{g_{0}},$ (32)
so that equations 30 yield a set of $3(k-1)$ RRs for them. For $k\leq 3$, and $z_{A}=z_{B}=z_{C}=z$, such RRs always converge to a fixed point of type $A_{i}=B_{i}=C_{i}$, corresponding to the isotropic phase. For $k\geq 4$, besides the isotropic fixed point, a nematic one is stable for large $z$, where, e.g., $A_{i}\neq B_{i}=C_{i}$.
Considering the simplified solution, the partition function of the system, obtained by attaching six sub-trees to the three vertices of the central triangle, reads:
$Y=s_{A}s_{B}s_{C}$ $+$ $z_{A}^{2}g_{0}^{2}s_{A}\sum_{n=0}^{k-2}g_{(n,A)}g_{(k-2-n,A)}+z_{B}^{2}g_{0}^{2}s_{B}\sum_{n=0}^{k-2}g_{(n,B)}g_{(k-2-n,B)}$ (33)
$+$ $z_{C}^{2}g_{0}^{2}s_{C}\sum_{n=0}^{k-2}g_{(n,C)}g_{(k-2-n,C)}=g_{0}^{6}y,$
where $y$ only depends on $z$ and on the ratios $A_{i}$, $B_{i}$ and $C_{i}$ at their fixed points, in the thermodynamic limit.
The density of monomers per site (at the vertices of the central triangle) belonging to rods in direction $j=A,B$ or $C$ is given by:
$\rho_{j}=\frac{z_{j}}{3y}\frac{\partial y}{\partial z_{j}},$ (34)
and, then, the total density is $\rho=\rho_{A}+\rho_{B}+\rho_{C}$. Following Ref. drs11 , the order parameter for the isotropic-nematic transition will be defined as
$\Psi=\frac{\rho_{A}-\rho_{B}}{\rho},$ (35)
assuming that $\rho_{B}=\rho_{C}$ always.
As demonstrated, e.g., in Ref. furlanALG , the bulk free energy per site, $\phi_{b}$, for the triangular HL is
$\phi_{b}=-\frac{k_{B}T}{3}\ln\left[\frac{Y^{\prime}}{Y^{4}}\right]=-k_{B}T\ln\left[\frac{f_{0}^{2}}{y}\right],$ (36)
so that
$\varphi_{b}=\frac{\phi_{b}}{k_{B}T}=-\ln\left[\frac{f_{0}^{2}}{y}\right],$ (37)
with $f_{0}$ being the convergent part of $g^{\prime}_{0}$, once we write $g^{\prime}_{0}=g_{0}^{4}f_{0}$.